Factorise 2Factorise+2 (a+b)^2} }}-9(a+b)-5?
Answer by stanbon(75887)

2(a+b)^2} -9(a+b) -5
----
Let a+b = w
---
Rewrite the problem:
2w^2 - 9w - 5
-----
= 2w^2-10w+w-5
----
= 2w(w-5)+(w-5)
----
= (w-5)(2w+1)
----
Change back to a+b
----
= (a+b-5)(2(a+b)+1)
---
= (a+b-5)(2a+2b+1)
=========================
Cheers,
Stan H.
=============